The Hiking Club plans to go camping in a
State park where the probability of rain on
any given day is 0.7. Which expression can
be used to find the probability that it will rain
on exactly three of the seven days they are
there?

Answer by htmentor(1343)

This is a binomial probability problem with the probability of "success"
equal to 0.7. The probability of x successes in n trials is given by:
P(x) = C(n,x)p^x(1-p)^(n-x), where C(n,x) is the number of combinations
(n things taken x at a time)
Thus P(3) = C(7,3)*0.7^3*0.3^4 = 0.0972